### v1

# Semantic Chunking and the Entropy of Natural Language

Abstract
The entropy rate of printed English is famously estimated to be about one bit per character, a benchmark that modern large language models (LLMs) have only recently approached. This entropy rate implies that English contains nearly 80 percent redundancy relative to the five bits per character expected for random text. We introduce a statistical model that attempts to capture the intricate multi-scale structure of natural language, providing a first-principles account of this redundancy level. Our model describes a procedure of self-similarly segmenting text into semantically coherent chunks down to the single-word level. The semantic structure of the text can then be hierarchically decomposed, allowing for analytical treatment. Numerical experiments with modern LLMs and open datasets suggest that our model quantitatively captures the structure of real texts at different levels of the semantic hierarchy. The entropy rate predicted by our model agrees with the estimated entropy rate of printed English. Moreover, our theory further reveals that the entropy rate of natural language is not fixed but should increase systematically with the semantic complexity of corpora, which are captured by the only free parameter in our model.
Figure 1: Schematic overview of two routes for estimating the entropy of a text.
(a) The same text is provided to an LLM either to compute token log-probabilities (b) or to perform semantic chunking (e). (c) The LLM perplexity (equivalently, the per-token cross-entropy loss on the input) is converted into an LLM-based estimate of the text’s entropy rate, $h_{\text{LLM}}$ (y-axis in (d)). (f) The chunking procedure applies recursive semantic segmentation until the single-token level, thereby parsing the document into a hierarchical tree of spans whose leaves are tokens (a “semantic tree”). (g) Example semantic tree for the input text. (h) The probability of observing the resulting tree structure is computed under a random-tree ensemble model. (i) Example combinatorial tree from the ensemble corresponding to the semantic tree in (g). (j) Ensemble probabilities are converted into a theoretical entropy-rate estimate $h_{\text{theory}}$, which closely matches $h_{\mathrm{LLM}}$ across diverse corpora ((d)).
I Introduction
Reading and understanding a natural text involves gradual comprehension of its meaning at different levels of abstraction, from single words to specific details, and finally to its overall gist. This hierarchical structure of texts generates strong statistical redundancies at all levels: words in a sentence have to satisfy the grammar, sentences have to relate to each other in a particular way, and the end of the story has to correspond to how the story begins. A reasonable, if necessarily simplified, way to describe the hierarchical structure of texts is by “semantic trees,” where each node corresponds to a particular “keypoint” that summarizes a certain part of the text [1, 2, 3, 4]. Every text will be characterized by its unique tree as its meaning develops, with different people forming different trees for the same text depending on how they comprehend it. In our recent publication [4], we argued that when considering multiple texts of a particular length, the corresponding ensemble of trees can be approximated by a random recursive partition of texts into shorter and shorter chunks. Moreover, we assumed that the splitting parameter of this ensemble, namely the maximal number of children for each node, is limited by working memory capacity.
The entropy of natural language is inherently related to its semantic structure. The redundancy of natural language implies that one can often anticipate an upcoming word from its context. Crucially, this anticipation is not a single-step guess but operates through a hierarchy of abstract inferences: a reader first infers the global topic and communicative intent, then the local discourse function of the current section and paragraph, and only subsequently progressively refines expectations down to the immediate grammatical structure that determines the next word [5, 6, 7, 8]. This nested cascade of inferences is mirrored in the structure of the text itself in the form of semantic trees. These semantic chunks encode information about possible continuations before any word- or token-level realization is specified. Accordingly, the information captured by the semantic structure constitutes a component of the total entropy and can be viewed as a lower bound on the full entropy rate of language.
On the other hand, the redundancy of language can be quantified by its entropy rate, relative to the maximal possible rate that depends on the size of the alphabet (i.e., the number of characters or words). The first estimate of the entropy rate of printed English was obtained by Shannon using the now-famous “guessing” game procedure, where a subject was trying to guess the next character after reading passages of text of various lengths [9, 10]. The resulting estimate of about 1 bit per character was then confirmed in other languages as well [11, 12]. Current large language models (LLMs) can now be used to estimate entropy rates of language without using human subjects and result in similar estimates [13, 14, 15, 16]. Until now, however, there is no first-principles understanding of what accounts for the entropy rate of English and other languages. The goals of this study are two-fold. First, we apply LLMs to segment multiple texts from several corpora, ranging from children’s books to modern poetry. We compare the obtained set of trees to a random ensemble proposed in [4] and find the splitting parameter that yields the best match in terms of chunk-size distribution. We then develop a theoretical framework for calculating the entropy of the random tree ensemble and compare the results with empirically obtained entropy rates across different corpora. Finally, we propose that the entropy rate of a particular text correlates with its comprehension difficulty and suggest future experiments to test this hypothesis.
II Entropy from LLM Perplexity
We define the entropy rate of language as the uncertainty associated with the next word/token given its preceding context. Auto-regressive large language models (LLMs) [17, 18] are trained to approximate this conditional distribution by minimizing cross-entropy loss on observed text [19]. At inference time, rather than using an LLM for generation, we can use its context window as a measurement device to estimate the entropy rate of a given sequence.
Concretely, for a text represented as a length-$N$ token sequence $\{t_{i}\}_{i=1}^{N}$, we evaluate the model left-to-right by conditioning on the prefix $t_{<i}$ and recording the probability $P(t_{i}|t_{<i})$ assigned to the observed token $t_{i}$ for each position $i\in\{2,\dots,N\}$. The resulting per-token surprisal sequence, $\log P(t_{i}|t_{<i})$, averaged over tokens, yields the model’s cross-entropy rate estimate. By standard information-theoretic arguments, this quantity upper-bounds the true entropy rate of the text distribution (see Fig. 1(a)–(c)) [10, 20],
$h_{LLM}=-\frac{1}{N}\sum_{i=1}^{N}\log P(t_{i}|t_{<i}),$ (1)
which is commonly interpreted as the model’s per-token cross-entropy rate on the text [21]. It is also referred to as the log-perplexity [22, 23], $h_{\mathrm{LLM}}=\log\mathrm{PPL}$.
III Entropy from Semantic Chunking
While cross-entropy (log-perplexity) provides an estimate of the entropy rate via next-token prediction, it does not, by itself, explain which aspects of semantic organization account for this uncertainty. In what follows, we operationalize the idea of estimating entropy from semantic trees by using an LLM to identify semantically coherent “chunks” recursively, thereby inducing a hierarchical decomposition of the text (see Fig. 1(e)–(g)).
Figure 2: Compare chunk size distributions between the random tree model and empirical semantic trees obtained via recursive semantic chunking.
(a) Empirical versus theoretical chunk-size distribution at an intermediate tree level ($L=7$) for 20 narratives from RedditStories. (b) Normalized empirical chunk-size distributions (pooling chunk statistics from 100 narratives) compared with the theoretical prediction $f_{L}$ at multiple levels $L$.
Semantic parsing has a long history in computational linguistics [24, 25, 26, 27]. Much of this literature focuses on sentence- and sub-sentence–level structure, where syntactic and grammatical constraints are explicit and can be evaluated against well-defined formal criteria. Above the sentence level, discourse parsing can be pursued using rhetorical theories [2], but these approaches often depend on expensive manual annotation and exhibit limited scalability [28, 29].
The recent emergence of large language models (LLMs) has renewed interest in document-level segmentation, particularly because many practical applications require processing texts that exceed a model’s context window [30, 31, 32]. This has catalyzed the development of text segmentation methods commonly referred to as chunking [33, 34]. In this setting, embedding-based and agentic semantic chunking methods have become popular alternatives to fixed-size chunking, which partitions text into constant-length windows and often breaks semantic contiguity. In contrast, semantic chunking seeks to segment a document into contiguous, semantically coherent units, such that each chunk is as self-contained as possible with respect to meaning and linguistic information [35, 36, 37].
In principle, many different structural priors could be used to parse a text into a hierarchy—for example, priors derived from discourse theory [38, 39]. Motivated by our earlier work on human narrative memory [4], we adopt $K$-ary trees as a structural prior: at each level of the hierarchy, the text is partitioned into at most $K$ chunks. In prior work, closely related tree-structured representations captured key aspects of how humans comprehend and recall narratives in large-scale memory studies [40].
We use an LLM to recursively identify semantically coherent “chunks” at multiple scales (see SI for the full algorithm). Starting from the full document, we partition the text into at most $K$ contiguous, semantically coherent chunks (Fig. 1(e)). We then apply the same procedure to each chunk recursively, continuing until reaching the single-token level (Fig. 1(f)). The result is a hierarchical tree of text spans, in which individual tokens appear as leaves (blue circles in Fig. 1(g)) and internal nodes correspond to coherent spans at progressively coarser resolutions (green circles in Fig. 1(g)).
We model the induced token tree as a random weak-integer ordered-partition process (Fig. 1(i); details in the next section). Over a large corpus, the collection of such trees forms an empirical approximation to a $K$-ary random-tree ensemble that admits analytic treatment. Using results from the theory of random $K$-ary trees, we compute the probability of each observed semantic tree associated with a given text (Fig. 1(h)). Converting these tree probabilities into Shannon information with respect to the ensemble yields an entropy estimate for the semantic structure (Fig. 1(j)), which we then compare against the LLM-based cross-entropy estimate. Surprisingly, the two entropy estimates agree closely across a diverse range of corpora, spanning children’s books, narrative fiction, arXiv abstracts, and modern poetry (Fig. 1(d)).
IV Semantic Trees
In this section, we first summarize the key theoretical results that characterize random $K$-ary trees. We then present empirical results from the semantic trees induced by our recursive chunking procedure, and show that random $K$-ary trees provide a good approximation to these semantic trees at the corpus level.
IV.0.1 Theory of Random Tree Ensemble
To quantify the empirical semantic trees, we associate each node (text span) with its token count (green circles in Fig. 1(g),(i)). These counts correspond to the lengths of semantically coherent units (e.g., paragraphs, sentences, or phrases). Across a large corpus, the resulting multiscale length statistics appear stochastic. To model this hierarchical representation, we adopt a self-similar splitting process [41, 42, 4] that reduces to a weak integer ordered partition problem: a text of $N$ tokens is divided into $K$ (possibly empty) chunks by placing K1K-1 boundaries uniformly at random between tokens (empty chunks are recorded as 0, as in Fig. 1(i)). We then apply the same procedure recursively to each non-empty chunk until all leaves have size 1. In principle, to avoid pathologies such as infinite recursion, we also allow non-unit leaves when a chunk fails to split at the level above (SI). In practice for the semantic trees , however, the vast majority of leaves are single tokens; leaves of size >1>1 typically correspond to multi-token words or expressions (e.g., idiomatic phrases) that the chunker tends to preserve as atomic units (SI). Conditioned on a parent chunk of size $n$, the size $m\geq 0$ of a given child is distributed as
$p_{\text{split}}(m|n)$ $=Z_{K-1}(n-m)/Z_{K}(n),$ (2)
$Z_{K}(n):$ $=\binom{n+K-1}{n-1}$, (3)
which defines an Markov chain over chunk sizes (SI).
Mathematically, this construction defines a one-parameter family (indexed by $K$) of random tree ensembles. A statistic of central interest is the distribution of chunk sizes at a fixed level $L$, conditioned on the root size $N$. We denote this distribution by $P_{L}(n|N)$ (and write $P_{L}(n)$ when $N$ is understood). Intuitively, $P_{L}(n)$ quantifies, for example, the probability that a node at level $L=3$ has size $n=4$.
We compute $P_{L}(n)$ by propagating the splitting kernel in Eq. (2) from the root to level $L$: we take products of the conditional probabilities across levels $l\leq L$ and marginalize over all intermediate node sizes while fixing the root to $N$ (see derivation in [4], also reproduced here in SI). The resulting theoretical distributions are typically concentrated near small $n$; as $N$ increases, probability mass shifts toward larger chunk sizes, reflecting the fact that longer texts admit larger spans at the same depth. Solid curves in Fig. 2(a) illustrate $P_{L}(n)$ for different values of $N$ at level $L=7$.
To test these theoretical predictions, we apply our semantic chunking algorithm to 100 stories randomly sampled from the RedditStories dataset [43], inducing one semantic tree per story. The algorithm has a single hyperparameter, the maximum branching factor $K$, which we set to $K=4$ (a choice we justify later in Sec. V.0.1). The sampled stories range from $\sim 100$ to $\sim 2500$ tokens. All resulting semantic trees have at least 7 levels of depth, and the longest stories reach depths of up to $\sim 12$ levels. Figure 2(a) compares the empirical chunk-size distributions from a subset of 20 semantic trees against the theoretical predictions (after re-normalizing the empty chunks), showing good agreement.
Because stories have heterogeneous lengths $N$, it is not a priori obvious how to aggregate statistics across texts. Next, we show that the theory admits a scaling limit for $N\gg 1$, enabling principled pooling and yielding corpus-level ensemble statistics for semantic trees.
IV.0.2 Large $N$ limit
For $N\gg 1$, the level-$L$ chunk-size distribution $P_{L}(n)$ admits a scaling form, where $P_{L}(n)$ converges to an $N$-independent continuous scaling function $f_{L}$ (parameterized by $L$ and $K$) [4]:
$P_{L}(n)$ $\simeq\frac{1}{N}f_{L}(n/N),$ (4)
$f_{L}(s)$ $=\begin{cases}\displaystyle\int_{s}^{1}\frac{dr}{r}\,p_{B}(r)\,f_{L-1}\!\left(\frac{s}{r}\right),&L>1,\\[9.47217pt] \displaystyle\delta(s-1),&L=1.\end{cases}$ (5)
That is, successive levels are related by a multiplicative convolution with a Beta-distributed random variable $r\in(0,1)$,
$p_{B}(r)=(K-1)(1-r)^{K-2}=\text{Beta}(1,K-1).$ (6)
Using the scaling function $f_{L}$, we can remove explicit $N$-dependence by pooling normalized chunk sizes $s_{L}:=n/N$ at a fixed level $L$ across all stories, and treating them as samples from a single empirical distribution $\hat{f}_{L}$. This pooling procedure enables a direct comparison between $\hat{f}_{L}$ and the theoretical scaling function $f_{L}$ using all chunks from the 100 semantic trees at each level (Fig. 2(b)). We find that $\hat{f}_{L}$ is well captured by $f_{L}$ across levels, with deviations emerging only at larger depths where finite-sample effects become appreciable (e.g., $L=11$). The match between theory and data indicates that semantic trees produced by recursive chunking lie in the random tree ensemble’s scaling regime.
Figure 3: Entropy rates across corpora.
(a) Entropy rate as a function of $K$. Theoretical values correspond to $h_{K}$ in Eq. (11). Empirical values are obtained by selecting the optimal $K$ for each corpus (Table 1) and comparing $h_{K}$ to the corpus-level LLM estimate $h_{\mathrm{LLM}}$. Shaded bands indicate three approximate entropy-rate regimes that differentiate genres. (b) Entropy-rate estimates from individual random-tree realizations (simulated using Eq. (2)) concentrate around the predicted value as $N$ increases, consistent with the emergence of typical trees. (c) Per-token information for 100 RedditStories computed in two ways: from LLM perplexity and from the likelihood of empirical semantic trees obtained via chunking. As $N$ increases, both estimates fluctuate around the predicted entropy rate. (d)–(f) Cumulative LLM surprisal $\sum_{i=1}^{N}\log P(t_{i}|t_{<i})$ for 100 texts from each corpus (color indicates text length $N$). The blue dash-dotted fit gives the LLM-based entropy-rate estimate $h_{\mathrm{LLM}}$. The red dashed line shows the theoretical prediction $h_{K}$ using the optimal $K$ for each corpus; its intercept is chosen for visual comparison (matched to the blue fit).
V From Semantic Trees to Entropy
Having established a corpus-level statistical description of the semantic trees, we now use the ensemble to compute an entropy associated with tree structure and compare it to $h_{\text{LLM}}$.
The model has a single parameter, $K$, which sets the maximum branching factor between successive levels. For fixed $K$, we represent a particular tree realization as $T=(\mu_{1},\dots,\mu_{L}):=(\mu_{l})_{l=1}^{L}$, where $\mu_{l;i}$ denotes the size of the $i^{\mathrm{th}}$ node at level $l$. Under the splitting process described above, the probability of a specific tree configuration factorizes as
$P(T)=\prod_{l=1}^{L}\prod_{i=1}^{K^{l-1}}Z_{K}(\mu_{l;i})^{-1}.$ (7)
This form follows from the fact that, conditioned on a parent node size, the $K$-tuple of child sizes is uniformly distributed over weak ordered partition (SI), and that there are $K^{l-1}$ nodes at level $l$ when zero-sized (empty) nodes are included. We emphasize, however, that conditioning on the root and marginalizing over intermediate levels induces highly non-uniform marginals such as $P_{L}(n)$, as seen in Fig. 2(a).
Table 1: Goodness-of-fit, measured by the average KL divergence $\langle D_{\mathrm{KL}}\rangle$, across corpora [44, 45, 43, 46, 47] for different values of $K$. For each corpus, the optimal $K^{\star}=\arg\min_{K}\langle D_{\mathrm{KL}}\rangle$ is highlighted in green.
$K$ TinyStories FairytaleQA RedditStories arXivAbstracts ModernPoetry
[…]
55 0.318 0.572 0.111 0.415 0.172
[…]
77 N/A N/A N/A N/A 0.174
88 N/A N/A N/A N/A 0.307
Using Eq. (7), we can write the Shannon entropy of the size-$N$ random-tree ensemble, i.e., the distribution over all tree configurations with root size $N$, directly in terms of the level-wise chunk-size distributions $P_{L}(n)$ (SI):
$H(N)$ $=-\sum_{\begin{subarray}{c}T\in\text{tree}\\ \text{of size }N\end{subarray}}P(T)\log P(T)$ (8)
$=\sum_{L=1}^{N}\langle\log Z_{K}(n)\rangle_{n\sim\rho_{L}(n)},$ (9)
where $\rho_{L}(n)$ is the density of states measuring the number of internal nodes of size $n$ at level $L$,
$\rho_{L}(n)$ $=\begin{cases}\displaystyle K^{L-1}P_{L}(n),&2\leq n\leq N-1\\[9.47217pt] \displaystyle 0,&\text{otherwise}.\end{cases}$ (10)
Equation (9) is exact and admits an intuitive interpretation. By the Markov property of the splitting process, the tree entropy decomposes additively across levels and, within each level, into contributions from individual nodes. The per-node contribution, $\log Z_{K}(n)$, is the log-multiplicity of possible $K$-way splitting configurations for an internal node of size $n$. The factor $\rho_{L}(n)$ weights this contribution by the number (multiplicity) of internal nodes of size $n$ present at level $L$.
The full mathematical analysis of the entropy of the above tree ensemble will be published in a separate publication [48]. There we will show that in the limit of large $N$, the entropy in Eq. (9) is asymptotically extensive in $N$:
$H(N)\simeq h_{K}N.$ (11)
The leading coefficient $h_{K}$ (the entropy rate) depends only on $K$ and will be analyzed in detail in [48]. Numerical solution of Equation (9) for the entropy rate as a function of $K$ is shown in (Fig. 3(a); see red line). In the SI, we derive an analytic expression of $h_{K}$, for $K=2$ and in the asymptotic limit $K\gg 1$.
Importantly, Eq. (11) is not merely an ensemble-average statement. As will be shown in [48], the ensemble of trees exhibits an Asymptotic Equipartition Property: namely, in the limit of large $N$ the normalized negative log-likelihood of a tree converges (in probability) to the entropy rate:
$-\frac{1}{N}\log P(T)\xrightarrow{N\to\infty}h_{K}$ (12)
i.e. for large $N$ most of the trees generated according to the above procedure belong to a set of “typical trees” with per-leaf log-likelihooeds concentrated around entropy rate; total probability of this set approaches $1$ in this limit. This concentration is illustrated in Fig. 3(b): simulations at increasing $N$ show that realization-to-realization fluctuations in the estimated entropy rate vanish, converging to the theoretical value $h_{K}\approx 2.5$ nats/leaf for $K=4$.
We can now estimate the entropy rate of a text in two complementary ways: (i) from the LLM perplexity (Eq. (1)); and (ii) from the semantic-tree likelihood under the random-tree ensemble, by computing $\frac{1}{N}\log P(T)$ for the empirical semantic tree $T$ produced by chunking with $K=4$ (Eq. (7); the choice of $K$ is justified in Sec. V.0.1). Figure 3(c) compares these two estimates for 100 texts from RedditStories, using the Llama family of models (semantic chunking with Llama-4-Maverick [49] and perplexity with Llama-3-70B [50]). For shorter texts (small $N$), the tree-based estimate (green crosses) lies systematically below the LLM cross-entropy estimate (green dots). As $N$ increases, this gap shrinks, and both estimates converge to fluctuations around the predicted entropy rate $h_{K}$ (red dashed line).
V.0.1 Selecting $K$ for each corpus
The random-tree ensemble has a single free parameter: the maximum branching factor $K$. Empirically, $K$ corresponds to the maximum number of non-empty chunks (i.e., the number of keypoints) permitted at a given level in the semantic tree induced by segmentation. In practice, $K$ imposes an implicit structural prior and controls the segmentation granularity. If $K$ is chosen too large, a human or LLM chunker may over-segment, effectively collapsing multiple hierarchical levels into a single step. Conversely, if $K$ is too small, the chunker may under-segment, causing spans to terminate prematurely as leaves. In principle, the effective $K$ could vary across texts and across segmentation procedures, and it may not be uniquely identifiable for a single document. To reduce this ambiguity, we adopt a corpus-level simplifying assumption: each corpus admits an optimal branching factor $K^{\star}$ that best describes the ensemble statistics of its semantic trees. Operationally, we select $K^{\star}$ by fitting the random-tree model to the corpus-derived semantic trees and choosing the value of $K$ that yields the best agreement.
Empirically, we apply the semantic chunking algorithm to 50 texts from each corpus, sweeping $K$ from $2$ to $8$. For each choice of $K$, we compute the empirical level-wise chunk-size distributions induced by chunking and compare them to the corresponding theoretical predictions at the same $K$ (as in Fig. 2(d)). To quantify goodness of fit, we measure the discrepancy between empirical and theoretical ensembles using the KL divergence, averaged across tree levels.
$\langle D_{\mathrm{KL}}(\text{data}\,\|\,\text{theory})\rangle=\frac{1}{L}\sum_{l=1}^{L}\sum_{s\in\mathcal{S}_{l}}\hat{f}_{l}(s)\log\frac{\hat{f}_{l}(s)}{f_{l}(s)},$ (13)
where $\mathcal{S}_{l}$ denotes the support (bins) of normalized chunk sizes at level $l$, and $\hat{f}_{l}$ and $f_{l}$ are the empirical and theoretical scaling functions, respectively. We find that some values of $K$ yield substantially better agreement than others; the resulting $\langle D_{\mathrm{KL}}\rangle$ values across corpora and $K$ are reported in Table 1. We define the optimal branching factor for a corpus as $K^{\star}=\arg\min_{K}\langle D_{\mathrm{KL}}\rangle$, corresponding to the highlighted entry in each column.
Figure 4: Comparison between the random tree model and empirical semantic trees obtained via recursive semantic chunking.
(a) Theoretical scaling functions $f_{L}$ across levels $L$ shown on log–log axes. (b) When replotted using the $O(1)$ lognormal variable $x=(\ln s-\mu_{L})/\sigma_{L}$, the curves in (a) collapse onto the universal $\mathcal{N}(0,1)$ distribution as $L$ increases. (c) Empirical chunk size distributions $\hat{f}_{L}$ (same as in Fig. 2(b), rebinned to be uniform in $\log s$) across levels $L$. (d) The empirical curves in (c) likewise collapse under the transformation $x=(\ln s-\hat{\mu}_{L})/\hat{\sigma}_{L}$, consistent with the theoretical prediction.
To estimate the corpus-level LLM entropy rate $h_{\mathrm{LLM}}$, we perform linear regression of the cumulative surprisal (negative log-probability) on text length for 100 texts from each corpus; the fitted slope gives the mean per-token cross-entropy rate, equivalent to the perplexity-based estimate in Eq. (1). Figures 3(d)–(f) show these cumulative-surprisal curves for three corpora: TinyStories, RedditStories, and ModernPoetry, together with the theoretical prediction $h_{K^{\star}}$ computed using the optimal $K^{\star}$ from Table 1. Although individual texts within a corpus vary (as reflected by the spread of curves), the mean entropy rates differ markedly across corpora: approximately $1.2$ nats/token for TinyStories [44], $2.5$ nats/token for RedditStories, and $3.2$ nats/token for ModernPoetry [47], with the highest about three times the lowest.
The corpus-level entropy-rate estimates from LLM perplexity, $h_{\mathrm{LLM}}$ (obtained as the regression slopes in Fig. 3(d)–(f)), are plotted against the corresponding optimal $K^{\star}$ values in Fig. 3(a). The theoretical curve $h_{K}$ evaluated at these same $K^{\star}$ values provides a parameter-free prediction for each corpus; these predictions are the quantities compared against $h_{\mathrm{LLM}}$ in Fig. 1(d). We find close agreement: across corpora, the measured $h_{\mathrm{LLM}}$ lies near the theoretical value $h_{K^{\star}}$ (Fig. 3(a)). The same consistency is visible within RedditStories (Fig. 3(c)), where both the LLM-based and tree-based estimates fluctuate around the predicted rate (red dashed line) as $N$ increases. However, we caution that this agreement should be interpreted at the corpus level. Because we assign a single $K^{\star}$ to an entire corpus, the tree-based entropy for an individual text is not expected to match the LLM cross-entropy exactly; rather, it matches on average, with substantial text-to-text variability that is not captured by the corpus-level model (Fig. 3(c)).
Furthermore, the inferred $K^{\star}$ aligns with intuitive notions of textual complexity (Fig. 3(a)). Simpler corpora such as children’s stories (red band; TinyStories and Fairytales) exhibit lower entropy rates and smaller $K^{\star}$ than narrative and expository corpora (green band; RedditStories and arXivAbstracts), which in turn fall below more stylistically atypical corpora such as poetry (blue band; ModernPoetry). Thus, both the entropy rate and the optimal branching factor suggest systematic differences across genres.
VI Scaling in $L$ and universality of trees
In this section, we discuss further theoretical developments of the random tree model and test it againt data. After taking the large-$N$ limit in Eq. (4), the normalized chunk-size distribution depends only on $L$ and $K$ through the continuous scaling function $f_{L}$. Using renormalization-group analysis (SI), we show that for $L\gg 1$, $f_{L}$ converges to a lognormal family whose parameters are determined by $K$ and $L$:
$f_{L}(s)$ $\simeq\log\mathcal{N}(\mu_{L},\sigma_{L}^{2}),$ (14)
$\mu_{L}$ $=-(L-1)H_{K-1},$ (15)
$\sigma^{2}_{L}$ $=(L-1)H^{(2)}_{K-1}$, (16)
where $H_{K-1}$ and $H^{(2)}_{K-1}$ denote the harmonic numbers of the first and second kind, respectively.
Equation (14) implies a universality statement: after transforming to the $O(1)$ lognormal variable
$x_{L}:=\frac{\ln s_{L}-\mu_{L}}{\sigma_{L}},$ (17)
the level-dependent scaling functions $f_{L}$ should collapse onto a universal standard normal distribution, $\mathcal{N}(0,1)$, independent of $L$ (and, after standardization, independent of $K$ as well). Consistent with this prediction, the functions $f_{L}$ differ substantially across levels when plotted in the original normalized-size coordinate $s_{L}$ (Fig. 2(e)). However, when replotted as functions of $x_{L}$, the same curves collapse onto $\mathcal{N}(0,1)$ as $L$ increases (Fig. 2(f)). We observe the same behavior empirically. The level-dependent empirical distributions in Fig. 2(g) collapse onto the same universal curve when expressed in the standardized coordinate $x_{L}=(\ln s_{L}-\hat{\mu}_{L})/\hat{\sigma}_{L}$, where $\hat{\mu}_{L}$ and $\hat{\sigma}_{L}$ are the empirical mean and standard deviation of $\ln s_{L}$ at level $L$ (Fig. 2(h)).
Discussion
Natural language is simultaneously structured and uncertain: readers experience texts as hierarchies of coherent semantic units, yet each word remains only partially predictable from context. On the one hand, linguistics and cognitive science have developed discourse-level theories in which meaning is scaffolded by hierarchical organization, such as coherence relations, intentional structure, and rhetorical segmentation, providing an explicit account of how larger spans constrain the interpretation of smaller ones [38, 2, 51]. In these frameworks, predictability is an emergent consequence of how discourse structure narrows the space of plausible continuations.
Modern NLP operationalizes predictability via per-token surprisal (cross-entropy, or log-perplexity) under an auto-regressive model. These information-theoretic estimates link linguistic uncertainty to redundancy and compression, and to behavioral signatures of processing difficulty such as reading time, eye movements, and readability [52, 53, 54, 55]. This perspective traces back to Shannon’s “guessing game” and has been refined in psycholinguistics.
In this work, we connect these two seemingly separate notions of redundancy in natural language through a minimal, first-principles theory that links multiscale chunking to the predictability of language. We find semantic structures by recursively segmenting texts into coherent chunks, yielding a semantic tree whose nodes represent text spans at multiple levels of granularity. We model these trees with a one-parameter random $K$-ary tree ensemble that is analytically tractable and predicts that level-wise chunk-size distributions are log-normally distributed. These signatures are observed in LLM-induced semantic trees across multiple open corpora.
The central finding is that semantic structure quantitatively predicts token-level entropy. The semantic-tree likelihood under the ensemble yields one estimate of entropy rate, while an auto-regressive LLM provides an independent estimate via token-level surprisals. Across corpora, the ensemble-based prediction closely matches the LLM measurements, implying that a substantial fraction of token-level unpredictability is already encoded in the multiscale semantic decomposition. In this way, our results reconcile two complementary views of language: as a probabilistic token sequence and as a hierarchical semantic object.
The random-tree ensemble was first developed in the context of human narrative memory, where $K$ plays the role of a working-memory capacity parameter [4], a natural psycholinguistic interpretation is that $K^{\star}$ here reflects working-memory load during comprehension, roughly, the number of concurrently active semantic chunks a reader must maintain to sustain coherent interpretation. Concretely, this could correspond to the number of keypoints needed to track a plot, or the number of interacting phrases required to resolve the meaning of a sentence. Notably, the empirically selected values $K^{\star}\in[2,6]$ fall within the range typically associated with human working-memory limits. From this perspective, the perceived complexity of poetry relative to children’s texts has a quantitative explanation: it induces a higher effective load on working memory during comprehension. This connection between entropy rate or information content of text to working memory remains one of the key future direction of investigations.
Acknowledgements.
W.Z. acknowledges funding from NSF ACCESS Discover Grant CIS240836, and support from the Simons foundation and the Eric and Wendy Schmidt Membership in Biology at the Insitute for Advanced Study. MK is supported in part by a grant from Fran Morris Rosman and Richard Rosman. M.T. is supported by the Simons Foundation, MBZUAI-WIS Joint Program for Artificial Intelligence Research and Foundation Adelis.
Supplementary Information for
“Semantic Chunking and the Entropy of Natural Language”
I Markov chain with absorbing states
Let’s consider a tree model with following absorbing rules: (1) if a node fails to split at some level, i.e., all its sibling nodes are zero, or (2) if the node has size less than or equal to 1, then it stops splitting further (this is the model described in the PRL paper).
To incorporate the absorbed states, for each node we can think of “doubling” the original $N+1$ states with its accompanying absorbed states: the first $N+1$ states 0,,N0,...,N are the absorbed states, which does not split in the next level; N+2,,2N+2N+2,...,2N+2 states are the survived states 0,,N0,...,N, where we identify the fictitious survived $0,1$ states as identical to the absorbed $0,1$ states. We can define the state vector as the probability distribution of chunk sizes at the $l^{th}$ level,
$\vec{P}_{L}:=\left(P(n^{(l)}=0^{\prime}),P(n^{(l)}=1^{\prime}),\dots,P(n^{(l)}=N^{\prime})|P(n^{(l)}=0),P(n^{(l)}=1),\dots,P(n^{(l)}=N)\right)^{\mathbf{T}}$ (S.1)
where we use prime to denote the absorbed states. Then we can write down the Markov transition
$\vec{P}_{L}=T^{L}\vec{P}_{0}$ (S.2)
with the transition matrix given by
$T\;=\;\left(\begin{array}[]{c|c}\mathbf{1}_{N+1}&\begin{array}[]{ccccc}1&0&0&\cdots&0\\ 0&1&0&\cdots&0\\ 0&0&\dfrac{1}{Z_{K}(2)}&\ddots&\vdots\\ \vdots&\vdots&\ddots&\ddots&0\\ 0&0&0&\cdots&\dfrac{1}{Z_{K}(N)}\end{array}\\ \hline\cr\mathbf{0}_{N+1}&\begin{array}[]{ccccc}0&0&*&\cdots&*\\ 0&0&*&\dfrac{Z_{K-1}(j^{\prime}-i^{\prime})}{Z_{K}(i^{\prime})}&\vdots\\ \vdots&\vdots&\ddots&\ddots&*\\ 0&0&\cdots&0&0\end{array}\end{array}\right)$ (S.3)
where $i^{\prime}=i-(N+1),j^{\prime}:=j-(N+1)$. The rows represent parent and columns represent their children, so all columns sum to one. The upper-left block simply means absorbed states stay absorbed. The lower-right block is our old upper-triangular Markov transition matrix with two new tweaks: (1) The first two columns are all zeros except for two entries (in the upper-right block) because we have identified the fictitious survived 0,1 states with the absorbed 0,1 states; (2) its diagonal are all zeros, because the diagonal entries represent the scenario where a node fails to split (size $n\to n$), therefore corresponds to a transition from survived state to absorbed state, and the previous diagonal entry $Z_{K}(i)^{-1}$ now simply gets moved to the diagonal in the upper-right block.
I.0.1 Occupation Number
Let’s call $T$ in Eq. (S.3) the node-size transition matrix. Instead of the node size distribution $\vec{P}_{L}$, we can also characterize the tree by its occupation number $\vec{m}_{L}$, where $(m_{L})_{n}:=$ number of nodes at level $L$ with size $n$. $\vec{m}_{L}$ is a $2(N+1)$ dimensional vector that characterizes the configurations of all the nodes at the same level $L$. For example, for $N=3$, $\vec{m}_{L}=(2,1,0,0,0,0,1,0)$ means that at level $L$, we have four nodes, three absorbed nodes with size 0, 0, 1, and one survived node with size 2. Note that the values sum up to 3. In general, normalization is enforced for the occupation number through
$N=\vec{\chi}\cdot\vec{m}_{L},$ (S.4)
where $\vec{\chi}=(0,1,...,N,0,1,...,N)$ is the node size vector.
The occupation number $\vec{m}_{L}$ is also a Markov chain,
$\vec{m}_{L}=Q^{L}\vec{m}_{0},$ (S.5)
where $\vec{m}_{0}=(0,...,0,1)$. The occupation transition matrix is related to the node-size transition matrix $T$ via a diagonal branching ratio matrix $B$,
$Q=TB;\;\;B_{ij}=\delta_{ij}b_{j},$ (S.6)
where
$b_{i}=\begin{cases}1,\quad i\in\text{absorbed};\\ K,\quad\text{otherwise}.\end{cases}$ (S.7)
The total number of nodes up to and including level $L$ is given by
$C_{L}$ $=\vec{1}\cdot\vec{m}_{L}$ (S.8)
where $\vec{1}$ denotes a vector of $2(N+1)$ 1’s (all the states in the system).
II Large-$N$ limit
For completeness, we reproduce the derivation first reported in [cite] for taking the large-$N$ limit.
In the limit of large $N$, and assuming $N\gg K$, we can use Stirling’s approximation to write
$\log Z_{K}(N)=\log\frac{N^{K-1}}{(K-1)!}+O\left(\log n\right).$ (S.9)
Then we can rewrite the conditional probability of observing $n^{(l+1)}$ stars in a bin of the $(l+1)^{th}$ level
$P(n^{(l+1)}|n^{(l)})=\frac{Z_{K-1}(n^{(l)}-n^{(l+1)})}{Z_{K}(n^{(l)})},$ (S.10)
in terms of the normalized compression ratios $s_{l}=n^{(l)}/N$:
$\rho(s_{l+1}|s_{l})\equiv NP(Ns^{(l+1)}|Ns^{(l)})=N\frac{Z_{K-1}(N(s_{l}-s_{l+1}))}{Z_{K}(Ns_{l})}\approx(K-1)\frac{1}{s_{l}}\left(1-\frac{s_{l+1}}{s_{l}}\right)^{K-2}.$ (S.11)
The chunk size distribution at level $L$ is given by marginalizing the intermediate chunks in the following product of conditional probabilities
$P(n^{(L)})$ $=\sum_{{\begin{subarray}{c}1<n^{(L-1)}\\ <\dots<n^{(1)}\end{subarray}}}\prod_{l=1}^{L-1}P(n^{(l+1)}|n^{(l)})P(n^{(1)})$ (S.12)
$=\sum_{n^{(L-1)}}P(n^{(L)}|n^{(L-1)})P(n^{(L-1)}),$ (S.13)
$P(n^{(1)})$ $=\delta_{n^{(1)},N}\;.$ (S.14)
The normalized chunk size distribution defined in Eq. (4) is given by
$f(s_{L})=\prod_{l=1}^{L-1}\int_{s_{l+1}}^{1}ds_{l}\rho(s_{l+1}|s_{l})\rho(s_{1}|1)$ (S.15)
$\rho(s_{l+1}|s_{l}):=\frac{K-1}{s_{l}}\left(1-\frac{s_{l+1}}{s_{l}}\right)^{K-2}.$ (S.16)
II.1 Scaling exponent for small $n$
The scaling function $f(s_{L})$ diverges as $s_{L}\to 0$, which allows us to simplify the nested integrals.
As $s_{L}\to 0$, each of the integral $\int_{s_{l+1}}^{1}ds_{l}/s_{l}(1-\frac{s_{l+1}}{s_{l}})^{K-2}$ is dominated by their lower limit, and becomes essentially $\int_{s_{l+1}}^{1}ds_{l}/s_{l}$, therefore, we have very slow divergence
$\lim_{s_{L}\to 0}f(s_{L})=\frac{(K-1)^{L-1}}{(L-2)!}\left[\ln\frac{1}{s_{L}}\right]^{L-2}.$ (S.17)
II.2 Exact results for $K=2$
Let us skip to the notation
$f_{l}(s)$ $=\int_{s}^{1}\rho(s|s^{\prime})f_{l-1}(s^{\prime})ds^{\prime}$ (S.18)
$=(k-1)\int_{s}^{1}dt\frac{(t-s)^{k-2}}{t^{k-1}}f_{l-1}(t)$ (S.19)
For $l=1$
$f_{1}(s)=(k-1)(1-s)^{k-2}$ (S.20)
For $k=2$, things simplify considerably.
$f_{1}(s)$ $=1$ (S.21)
$f_{l}(s)$ $=\int_{s}^{1}dt\frac{1}{t}f_{l-1}(t)$ (S.22)
Seeing the first few cases instantly reveals the pattern
$f_{2}(s)$ $=\int_{s}^{1}\frac{dt}{t}=-\log(s)$ (S.23)
$f_{3}(s)$ $=\int_{s}^{1}\frac{dt}{t}\left(-\log(t)\right)=\int_{s}^{1}dt\partial_{t}\left(-\frac{1}{2}\log^{2}t\right)=\frac{1}{2}\log^{2}t$ (S.24)
Assume the ansatz
$f_{l}(s)=c_{l}\left(-\log s\right)^{l-1}$ (S.25)
Then
$f_{l+1}(s)$ $=c_{l}(-1)^{l-1}\int_{s}^{1}dt\partial_{t}\left(\frac{1}{l}\log^{l}t\right)=\frac{c_{l}(-1)^{l}}{l}\log^{l}(s)=c_{l+1}\left(-\log s\right)^{l}$ (S.26)
$\Rightarrow$ $c_{l+1}=\frac{c_{l}}{l}=\frac{c_{l-1}}{l(l-1)}=\frac{1}{\Gamma(l+1)}$ (S.27)
So that $c_{l}=1/\Gamma(l)$, and
$f_{l}(s)=\frac{1}{\Gamma(l)}\left(-\log s\right)^{l-1}$ (S.28)
The first moment
$\bar{s}=\int sf_{l}(s)ds=2^{-l}$ (S.29)
III Entropy of Trees
III.1 Proof of Eq. (9)
Given a tree configuration $(\mu_{1},\dots,\mu_{l})$, we simplify the notation by denoting $\mu_{l;i}$ as the size of the $i^{th}$ node at level $l$. The probability of $(\mu_{1},\dots,\mu_{l})$ is given by the product of the probabilities $q_{l-1}$ at each preceding level l1l-1,
$P(\mu_{1},\dots,\mu_{l}):$ $=\prod_{m=0}^{l-1}q_{m};$ (S.30)
$q_{m}:=\frac{1}{Z_{K}(\mu_{m})}$ $,\;\;\;q_{0}=\frac{1}{Z_{K}(N)}.$ (S.31)
At each level $l+1$, $q_{l}$ describes the uniform probability across all the children nodes splitted from their parents at level $l$. Therefore, $q_{l}$ is given by the product of the probabilities across all the nodes at level $l$
$q_{l}=\prod_{i=1}^{K^{l}}\frac{1}{Z_{K}(\mu_{l;i})}.$ (S.32)
The Shannon entropy at level $l+1$ is given by
$H_{l+1}$ $=-\sum_{\mu_{1}\dots\mu_{l+1}}P(\mu_{1},\dots,\mu_{l+1})\log P(\mu_{1},\dots,\mu_{l+1})$ (S.33)
$=-\sum_{\mu_{1}\dots\mu_{l+1}}q_{0}\cdots q_{l}\log(q_{0}\cdots q_{l})$ (S.34)
$=-\sum_{\mu_{1}\dots\mu_{l}}\left(q_{0}\cdots q_{l-1}\right)\left(\sum_{\mu_{l+1}}q_{l}\right)\log\bigg((q_{0}\cdots q_{l-1})q_{l}\bigg)$ (S.35)
where we note that the first parenthesis is $P(\mu_{1},\dots,\mu_{l})$ and second parenthesis equals one by definition of $q_{l}$. We can rewrite
$H_{l+1}$ $=-\sum_{\mu_{1}\dots\mu_{l}}P(\mu_{1},\dots,\mu_{l})\log P(\mu_{1},\dots,\mu_{l})-\sum_{\mu_{1}\dots\mu_{l}}P(\mu_{1},\dots,\mu_{l})\log q_{l}$ (S.36)
$=H_{l}+\Delta H_{l+1}.$ (S.37)
Using Eq. (S.32),
$\Delta H_{l+1}$ $=-\sum_{\mu_{1}\dots\mu_{l}}P(\mu_{1},\dots,\mu_{l})\log q_{l}$ (S.38)
$=\sum_{\mu_{1}\dots\mu_{l}}P(\mu_{1},\dots,\mu_{l})\sum_{i=1}^{K^{l}}\log Z_{K}(\mu_{l;i}),$ (S.39)
For reasons that will become clear later, let’s change the name of $\mu_{l;i}$ by introducing a new variable $n_{l}$ using $\sum_{n_{l}}\delta({n_{l}-\mu_{l;i}})$. Then
$\Delta H_{l+1}$ $=\sum_{n_{l}}\sum_{\mu_{1}\dots\mu_{l}}P(\mu_{1},\dots,\mu_{l})\sum_{i=1}^{K^{l}}\log Z_{K}(n_{l})\delta({n_{l}-\mu_{l;i}})$ (S.40)
$=K^{l}\sum_{n_{l}}\bigg[\sum_{\mu_{1}\dots\mu_{l}}P(\mu_{1},\dots,\mu_{l})\bigg(\frac{1}{K^{l}}\sum_{i=1}^{K^{l}}\delta({n_{l}-\mu_{l;i}})\bigg)\bigg]\log Z_{K}(n_{l})$ (S.41)
$=K^{l}\sum_{n_{l}}\left\langle\frac{1}{K^{l}}\sum_{i=1}^{K^{l}}\delta({n_{l}-\mu_{l;i}})\right\rangle_{\begin{subarray}{c}P(\mu_{1},\dots,\mu_{l})\end{subarray}}\log Z_{K}(n_{l})$ (S.42)
$=K^{l}\sum_{n_{l}}\left\langle\langle\delta(n_{l}-\mu_{l})\rangle_{\begin{subarray}{c}\text{bins at}\\ \text{level }l\end{subarray}}\right\rangle_{\begin{subarray}{c}\text{trees up}\\ \text{to level }l\end{subarray}}\log Z_{K}(n_{l})$ (S.43)
$=K^{l}\sum_{n_{l}}P_{l}(n_{l})\log Z_{K}(n_{l})$ (S.44)
$=K^{l}\sum_{n}\langle\log Z_{K}(n)\rangle_{P_{l}(n)}.$ (S.45)
where we have used the definition of the marginal distributions of node size at level $l$.
Since entropy comes from just the survived states, we define as in Eq. (9) $\rho_{L}(n):=K^{L}P_{L}(n)$ for $2\leq n\leq N-1$, we can simply write
$H(N)=\sum_{L=1}^{N}K^{L}\sum_{n=2}^{N-1}P_{L}(n)\log Z_{K}(n).$ (S.46)
Here we enumerate all the trees for a few cases by hand, and computed their entropy, which is reported below (the notation $(n_{1},...,n_{m})^{r}$ denotes the partition $(n_{1},...,n_{m})$ with $r$ permutations):
  • $N=3$, $K=2$. A total of 8 trees: $(3,0)^{2}$ with $p=1/4$; $\big((2,0)^{2},1\big)^{2}$ with $p=1/12$; $\big((1,1),1\big)^{2}$ with $p=1/12$. Entropy is $2\log 2+1/2\log 3\approx 1.936\text{ nats}$.
  • $N=4$, $K=3$. A total of 369 trees. Entropy is 5.134 nats.
  • $N=5,K=3$. A total of 3885 trees. Entropy is 6.983 nats.
These all agree with the numerical values computed from Eq. (S.46).
I also wrote a program to enumerate the trees and compare the entropy with Eq. (S.46), and the result agrees almost exactly, see Fig. S1(a). The curve is computed for $K=4$, it has some initial nonlinearity (for $N<5$) but soon becomes linear, with slope about 2.5 nats.
Figure S1: Entropy of trees. (a) Both theory and enumeration are for $K=4$. $H(N)\approx 2.5\text{ nats}\times N$. (b)-(c) Cumulative log probabilities (TI) for different stories (shown in different shades of green), blue solid curve corresponds to linear regression line, red dashed line corresponds to our theory computed for $K=4$ from Eq. (S.46). (b) 26 Labov stories taken from [40, 56, 4]. (c) 1000 RedditStories.
Interestingly, for both the Labov stories and 1000 randomly chosen RedditStories, the entropy rate is also about the same, 2.5 nats/token as predicted by theoretical entropy rate with $K=4$, see Fig. S1(b-c).
III.2 Entropy scaling
We first rewrite the full entropy formula Eq. (S.46) into a recursive one, which is useful for calculating the entropy numerically.
$H(N)=\ln Z_{K}(N)+K\sum_{n=2}^{N-1}\frac{Z_{K-1}(N-n)}{Z_{K}(N)}H(n).$ (S.47)
Below, in Sec. III.2.1 we first demonstrate a simple derivation of the entropy rate for the case of $K=2$ as an illustration. In Sec. III.2.2, we show that in the limit of large $K$ it is possible to prove that $H(N)$ scales linearly in $N$ (Eq. (11)) and obtain close-form solution for $h_{K}$.
III.2.1 Exact entropy rate for $K=2$
For the specific case of $K=2$, the partition functions simplify to $Z_{2}(N)=N+1$ and $Z_{1}(N-n)=1$. Substituting these into the general recursion Eq. (S.47), we obtain:
$H(N)=\ln(N+1)+\frac{2}{N+1}\sum_{n=2}^{N-1}H(n).$ (S.48)
This recurrence relation has the same form as the one governing the average number of comparisons in the ’quicksort’ array sorting algorithm, which can be solved exactly using a differencing trick, as shown by Knuth [57]. Cover and Thomas were the first to apply the same technique to calculate the entropy of random binary trees, given as an exercise in their seminal book on information theory (Problem 4.5 in [58]). While their exercise corresponds to random strong integer ordered partition with $K=2$, here we apply it to our weak ordered partition random tree model with $K=2$. Despite this difference in the underlying combinatorics, the form of the recurrence remains the same, and thus can be solved using the same differencing trick.
The first step is to get rid of the summation term by considering the difference (N+1)H(N)NH(N1)(N+1)H(N)-NH(N-1), which leads to an exactly solvable first-order linear recurrence relation. The solution is given by:
$H(N)=\ln(N+1)+2(N+2)\sum_{m=2}^{N-1}\frac{\ln(m+1)}{(m+2)(m+3)}.$ (S.49)
In the limit of large $N$, the entropy scales linearly as $H(N)\sim h_{2}N$. Extracting the coefficient of the linear term, we find the exact entropy rate for $K=2$:
$h_{2}=2\sum_{m=2}^{\infty}\frac{\ln(m+1)}{(m+2)(m+3)}.$ (S.50)
III.2.2 Large-$K$ limit, Mellin transform, and contour integral
Recall that the normalized chunk size distribution can be written as a Mellin convolution with beta-variables, Eq. (S.99), we can formally solve it by first applying the Mellin transform, then perform the inverse Mellin transform using contour integral. However, as will be noted at the end of this section, this argument only works for $K\gg 1$.
For a single Beta-variable $r\sim p_{B}(r)$, the Mellin transform of its pdf is
$\mathcal{M}[p_{B}](z)$ $=\int r^{z-1}p_{B}(r)dr$ (S.51)
$=(K-1)B(z,K-1),$ (S.52)
where $B$ is the Euler Beta function. Our distribution $f_{L}(s)$ describes the product of L1L-1 i.i.d. random beta variables (Eq. (S.100)), therefore the pdf in the Mellin space multiplies, and we have
$\mathcal{M}[f_{L}](z)$ $=\left(\mathcal{M}[p_{B}]\right)^{L-1}$ (S.53)
$=(K-1)^{L-1}B(z,K-1)^{L-1},$ (S.54)
which also means the moments of $f_{L}$ are (by definition of Mellin transform)
$m_{t}=(K-1)^{L-1}B(t+1,K-1)^{L-1}.$ (S.55)
Applying the inverse Mellin transform to Eq. (S.53), we have the spectral representation of the normalized chunk size distribution
$f_{L}(s)=\frac{1}{2\pi i}\int_{c-i\infty}^{c+i\infty}dzs^{-z}(K-1)^{L-1}B(z,K-1)^{L-1},$ (S.56)
where $c$ is taken to be some positive constant.
Note that when taking the large $N$ limit, the chunk size distribution $P_{L}(n)$, which is a discrete probability mass function, cannot converge point-wise to the p.d.f. $f_{L}(s)$ because the support is different. To enforce normalization we assume the bin-wise convergence:
$P_{L}(n)$ $=\int^{n}_{n-1}\frac{1}{N}f_{L}\left(\frac{n^{\prime}}{N}\right)dn^{\prime}$ (S.57)
$=\int^{n/N}_{(n-1)/N}f_{L}(s)ds.$ (S.58)
Then in the large-$N$ limit, the entropy Eq. (S.46) becomes
$H(N)$ $=\frac{1}{N}\sum_{n=2}^{N-1}\ln Z_{K}(n)g\left(\frac{n}{N}\right),$ (S.59)
$g\left(\frac{n}{N}\right)$ $=N\sum_{L=1}^{N}K^{L-1}\int^{n/N}_{(n-1)/N}f_{L}(s)ds.$ (S.60)
Using Eq. (S.56), Eq. (S.60) becomes
$g\left(\frac{n}{N}\right)$ $=\frac{1}{2\pi i}\int^{n/N}_{(n-1)/N}ds\int_{c-i\infty}^{c+i\infty}dzs^{-z}\sum_{L=1}^{N}\left[K(K-1)B(z,K-1)\right]^{L-1}$ (S.61)
$=\frac{1}{2\pi i}\int^{n/N}_{(n-1)/N}ds\int_{c-i\infty}^{c+i\infty}dz\frac{s^{-z}}{1-K(K-1)B(z,K-1)}.$ (S.62)
The integrand
$F(z):=\frac{s^{-z}}{1-K(K-1)B(z,K-1)}$ (S.63)
takes the form
$F(z)$ $=\frac{s^{-z}}{h(z)};$ (S.64)
$h(z)$ $:=1-K(K-1)B(z,K-1).$ (S.65)
The poles in $F$ corresponds to the zeros $z_{j}\in\mathbb{R}$ of $h(z)$
$1=K(K-1)B(z_{j},K-1),$ (S.66)
which exists in the complex plane leftward of $z=2$, and $\text{max}_{j}\text{Re}(z_{j})=2$. We can perform a Laurent expansion of the denominator $h(z)$ around the zeros
$h(z)$ $=h^{\prime}(z_{j})(z-z_{j})+O((z-z_{j})^{2});$ (S.67)
$h^{\prime}(z)$ $=-K(K-1)B(z,K-1)\left[\psi(z)-\psi(z+K-1)\right],$ (S.68)
where $\psi(x):=d\ln\Gamma(x)/dx$ is the digamma function. In particular, at the zeros $h(z_{j})=0$, the derivative simplifies to
$h^{\prime}(z_{j})=\psi(z_{j}+K-1)-\psi(z_{j}).$ (S.69)
For integer $n_{j}$, it further simplifies into Harmonic numbers
$h^{\prime}(n_{j})$ $=H_{n_{j}+K-2}-H_{n_{j}-1},$ (S.70)
$H_{K}$ $=\sum_{i=1}^{K}\frac{1}{i}.$ (S.71)
All the poles are toward the left of $z=2$, some are complex some are real, depending on choice of $K$. Eq. (S.62) is a Bromwich type integral, and we choose $c>2$ and close the contour counterclockwise to the left with radius $R\to\infty$, where the integrand decays exponentially like $e^{\ln sR}$ ($\ln s\leq 0$). The contour therefore encloses all the poles of $F(z)$. By the Residue Theorem, we can write the integral Eq. (S.62) as
$g\left(\frac{n}{N}\right)$ $=\int^{n/N}_{(n-1)/N}ds\sum_{k}\text{Res}(F,z_{k}).$ (S.72)
The residue at the dominant pole ($z=2$) becomes
$\int^{n/N}_{(n-1)/N}ds\text{Res}(F,z=2)$ (S.73)
$=$ $\frac{N}{H_{K}-1}\int^{n/N}_{(n-1)/N}s^{-2}ds$ (S.74)
$=$ $\frac{N^{2}}{(H_{K}-1)}\frac{1}{n(n-1)}.$ (S.75)
All other poles are sub-leading in $1/N$, $\text{Re}(z_{j})<2$, and their contributions $s^{-z_{j}}=(N/n)^{z_{j}}$ are exponentially smaller in $N$. All the poles are simple poles, because $h(z_{j})=0$ but $h^{\prime}(z_{j})\neq 0$. We can in fact calculate them in closed form. Let’s rewrite Eq. (S.68) using Gamma function recurrence relations
$h(z)$ $=1-\frac{K!}{Q(z)},$ (S.76)
$h^{\prime}(z)$ $=K!\frac{\sum_{m=0}^{K-2}(z+m)^{-1}}{Q(z)}$ (S.77)
$Q(z)$ $=\prod_{m=0}^{K-2}(z+m).$ (S.78)
Then the poles are simply the K1K-1 complex roots of the polynomial equation
$K!=Q(z).$ (S.79)
At these roots, the derivative simplifies to
$h^{\prime}(z_{j})=\sum_{m=0}^{K-2}\frac{1}{z_{j}+m}.$ (S.80)
Therefore, the residues are
$\text{Res}(F,z_{j})=\frac{s^{-z_{j}}}{\sum_{m=0}^{K-2}(z_{j}+m)^{-1}}.$ (S.81)
Then we have the asymptotic form of entropy
$H(N)$ $=\frac{1}{N}\sum_{n=2}^{N-1}\ln Z_{K}(n)g\left(\frac{n}{N}\right),$ (S.82)
$g\left(\frac{n}{N}\right)$ $=\int^{n/N}_{(n-1)/N}ds\sum_{j=0}^{K-1}\frac{s^{-z_{j}}}{\sum_{m=0}^{K-2}(z_{j}+m)^{-1}},$ (S.83)
$z_{j}$  $\in\text{ roots of }0=\prod_{m=0}^{K-2}(z+m)-K!.$ (S.84)
For $N\gg 1$, the dominant contribution is $z=2$. Going back to Eq. (S.59) we have
$H(N)$ $=h_{K}N+O(\log N),$ (S.85)
$h_{K}$ $=\frac{1}{H_{K}-1}\sum_{n=2}^{N-1}\frac{\ln Z_{K}(n)}{n(n-1)}.$ (S.86)
Recall that when deriving the recurrence relation for the pdf $f_{L}(s)$ in Eq. (S.94), we used Stirling approximation for $n\gg 1$ (Eq. (S.10)):
$Z_{K}(n)$ $=\binom{n+K-1}{K-1}$ (S.87)
$\approx\frac{n^{K-1}}{(K-1)!}.$ (S.88)
This approximation could fail in $Z_{K-1}(n_{L-1}-n_{L})$ when $n_{L-1}\sim n_{L}$. However, it becomes accurate when $K\gg 1$, because the typical size of $n_{L}$ is
$\langle n_{L}\rangle\sim\frac{n_{L-1}}{K}.$ (S.89)
Therefore, for large $K$, $n_{L}\ll n_{L-1}$ for most of the configurations. For small $K$, however, the edge contribution $n_{L-1}\sim n_{L}$ become non-negligible, and the error builds up. Away from the edge, the error is
$P(n|N)=\frac{K-1}{N}\left(1-\frac{n}{N}\right)^{K-2}+O\left(\frac{1}{N^{2}}\right).$ (S.90)
and subsequently
$P_{L}(n)=\int^{n/N}_{(n-1)/N}f_{L}(s)ds+O\left(\frac{L}{N^{2}}\right),$ (S.91)
which carries to the entropy
$H(N)=h_{K}N+O(\log N).$ (S.92)
To see how good the approximation is, we expand Eq. (S.86) for large $K$ ($\gamma\approx 0.577$ is Euler’s constant)
$h_{K}=\frac{1}{2}(\ln K)^{2}+(1+\gamma)\ln K+\frac{\pi^{2}}{12}-\ln 2+O(\frac{\ln K}{K}).$ (S.93)
In Fig. S2 we plot the numerics against Eq. (S.93), it appears to be an good approximation for $K\gg 1$.
Figure S2: Entropy scaling. Numerics is computed with the Markov model, large $K$ expansion is Eq. (S.93).
IV Recursive semantic chunking
Before chunking, we sanitize the user text by stripping an outer triple-backtick fence (if present), normalizing line endings, and converting tabs/newlines to single spaces. We then enforce sentence spacing by inserting a space after ”.”, ”!” or ”?” when the next character appears to start a new sentence (capital letter or digit), collapse multiple spaces into one, and trim leading/trailing whitespace. Finally, we remove control characters and delete all single/double quotation marks (ASCII and common Unicode variants), producing a stable, quote-free string whose remaining punctuation and interior spacing are consistent for both the chunker and verifier.
Then the text is passed into the chunking prompt (see below), and input to a LLM (LLama-4-Maverick) for segmentation. After receiving results returned from the chunker, they are pass into a verifier, where the verifier concatenate the chunks and check whether it equals the original text before chunking.
We use three chunking modes to balance minimal assumptions, robustness, and token efficiency across regimes. For most inputs, we default to sentence-level segmentation because it makes the weakest structural assumptions while still producing semantically coherent spans. For long texts, however, LLMs occasionally skip or paraphrase content, which breaks exact-match verification; to reduce this failure mode and to save input budget, we switch to a paragraph-level cut-point mode where the model only returns sentence cut point indices, and we only trigger this once the span exceeds a 200-token threshold, which is well above single-sentence level, so we are not imposing fine-grained sentence structure at that level. Finally, for very short spans (fewer than 6 tokens), models often resist splitting within words and may stop before reaching token granularity; in these cases, we provide an explicit token-indexed list and use a phrase-level cut-point prompt so the model can operate directly on token boundaries.
Below, we show the three different chunking prompts we use at the different levels:
Prompt 1 Text Segmentation (main)
System Prompt
You are a text-segmentation assistant. Split the entire input into up to *{K}* contiguous, non-overlapping segments such that concatenating them in order reproduces the original text exactly. Including ALL whitespaces and punctuation. Each segment should be semantically coherent. In the case of simple phrases, segment into semantically contiguous chunks of tokens. *CRITICAL*: Preserve ALL whitespace and quotation marks exactly as it appears in the input. If the input starts with a space, the first segment should start with that space. Return a raw JSON array of strings - do NOT wrap in markdown or code fences.
Examples
Input: 1"The quick brown fox jumps over the lazy dog." Output: 1["The quick brown fox", " jumps over", " the lazy dog."] Input: 1" brown fox" Output: 1[" brown", " fox"] Input: 1"the lazy dog." Output: 1["the", " lazy dog", "."] Input: 1" jumps over" Output: 1[" jumps", " over"]
User Prompt
Input text to segment: 1{text}
Prompt 2 Paragraph Segmentation (Cut-Point Mode)
System Prompt
You are a text-segmentation assistant **in CUT-POINT mode**. The input will be N numbered sentences (0 … N-1). Choose up to *{K-1}* cut points so that the resulting segments are semantically contiguous and most conceptually disjoint from the other segments. Return **raw JSON** -- a strictly ascending list of integers in the range 1 … N-1. Do NOT wrap in markdown.
Example
Input: 1[0] Yeah I was in the boy scouts at the time. 2[1] And we was doing the 50-yard dash racing but we was at the pier marked off and so we was doing the 50-yard dash. 3[2] There was about 8 or 9 of us you know, going down, coming back. 4[3] And going down the third time I caught cramps and I started yelling Help!’ but the fellows didnt believe me you know. 5[4] They thought I was just trying to catch up because I was going on or slowing down. Output: 1[1, 3]
User Prompt
Input numbered sentences: 1{sentences}
Prompt 3 Phrase-Level Segmentation (Token Indices)
System Prompt
You are a **phrase-level segmentation assistant**. Given a token list with indices, return a JSON array of cut-points (0-based) where a new segment starts. Choose no more than *{K-1}*, yielding up to *{K}* segments in total. Segments should be the largest contiguous group of tokens that form a coherent phrase (e.g. determiner + noun, adjective + noun, verb + particle). Attach sentence-final punctuation to the preceding word. Cut-points must be strictly increasing and between 1 and len(tokens)-1. **JSON only** -- respond with the RAW JSON array (e.g. [3]) and nothing else.
Example
Input: 1{"tokens": ["The", " quick", " brown", " fox"]} Output: Implied Segments (for illustration): 1["The quick brown", " fox"]
User Prompt
Input token list: 1{tokens}
IV.1 Examples semantic trees
In this section, we give a few examples of the resulting semantic trees from the algorithm described above.
Example Reddit Story (32721)
Due to the space constraints, we present a shorter story below (248 tokens), while a typical story are much longer. The following is an example (32721) completed by a Reddit user in response to another user-written prompt. The whitespaces and linebreaks have been normalized for display in below. The corresponding semantic tree for $K=4$ is shown in Fig. S3.
Prompt
“You have been enjoying Life in Heaven for over 100 Years and then suddenly, you wake up on a hospital bed …”
Story Excerpt
They depict heaven in all white in the books; the clouds, the Greek architecture, the angels in robes. White is the color of unity, the combination of all the colors of the visible spectrum. Heaven was white too, in a sense.
I woke up to a different type of white.
You’re back. Well of course I am. Hugs all around. It was pleasant, reflecting a bit of heaven on Earth. That’s what heaven was. Love. A feeling of completeness. Love makes us complete, that’s what it is. We were all pieces of the puzzle to a greater whole.
I can feel the warmth of everyone in the room. Yet I still feel the emptiness. Well of course I do. We humans can only convey a small portion of our feelings to one another. It was different in heaven. There was no you or I. Just us. Sharing is caring they say. Many who say it don’t realize the strength that the quote carries.
My wife rushes into the room in her work clothes. Eyes drenched with tears. She hugs me and I can simultaneously feel the warmth of her body and the wetness of the tears.
I’m glad you’re still here.
Figure S3: Semantic tree of Reddit story (32721). The tree was originally plotted as a long vertical figure and is adapted here into five panels. Read the tree from top to bottom and left to right across panels (a)–(e). The root appears in panel (c) as the green node (248).
IV.2 Example Tiny Story (198810)
In the following we give an example from the Tiny stories (198810), it has a total of 135 tokens. The corresponding semantic tree for $K=2$ is shown in Fig. S3.
Once there was a little girl who found a hidden band. She was so excited and took it with her to show her friends. When she showed it to her friends, they all wanted to play with it. But when the band got too close to the sun, it started to melt! All of the kids were so sad because the band was gone. But then one of them had an idea. They used their markers and colored paper to make a new band. It wasn’t as nice as the first, but it was still fun to play with! The little girl was so glad that they found a way to make a new band and they still had lots of fun!
Figure S4: Semantic tree of Tiny Story (198810). The tree was originally plotted as a long vertical figure and is adapted here into three panels. Read the tree from top to bottom and left to right across panels (a)–(c). The root appears in panel (b) as the green node (135).
IV.3 Example Modern Poetry (8443)
In the following we give an example from the Modern Poetries (8443), it has a total of 126 tokens. The corresponding semantic tree for $K=6$ is shown in Fig. S5.
Aiee! It is the ceremony of the first blades of winter. Horticulture, horticulture, the little steam train says puffing up the mountainside. As if he had never known a home of his own, only ditches. Three stomps with a stone stump and the colloquium started. Beggars under the drainpipe, another hand’s cast of the bone dice. Whatever name the event has, it can be understood as an invitation. Epilepsy, epilepsy, the little steam train said, descending at evening. They bowed so low that their wigs tangled and I had to laugh.
Figure S5: Semantic tree of Modern Poetry (8443). The tree was originally plotted as a long vertical figure and is adapted here into three panels. Read the tree from top to bottom and left to right across panels (a)–(c). The root appears in panel (b) as the green node (135).
V Renormalization of chunk size distribution
After taking the large $N$ limit, the normalized chunk size distribution is
$f_{L}(s)$ $=\int_{s}^{1}\rho(s|s^{\prime})f_{L-1}(s^{\prime})ds^{\prime}$ (S.94)
$=(K-1)\int_{s}^{1}\frac{ds^{\prime}}{s^{\prime}}(1-\frac{s}{s^{\prime}})^{K-2}f_{L-1}(s^{\prime})ds^{\prime}.$ (S.95)
Let
$r=\frac{s}{s^{\prime}},\qquad ds^{\prime}=-\frac{s^{\prime}}{r}dr,$ (S.96)
then
$f_{L}(s)=(K-1)\int_{s}^{1}\frac{dr}{r}(1-r)^{K-2}f_{L-1}(\frac{s}{r}).$ (S.97)
We can identify the Beta-distribution
$p_{B}(r)=(K-1)(1-r)^{K-2}=\text{Beta}(1,K-1),$ (S.98)
then
$f_{L}(s)=\int_{s}^{1}\frac{dr}{r}p_{B}(r)f_{L-1}(\frac{s}{r}),$ (S.99)
takes the form of a Mellin convolution, which describes the pdf of a product of two random variables. $f_{L}=p_{B}*f_{L-1}$. Therefore, the random variable for level-$L$ is a product of (L1)(L-1) i.i.d. Beta-random variables
$s=r_{1}r_{2}\cdots r_{L-1},\;\;r_{i}\sim p_{B}.$ (S.100)
For $L\to\infty$, $s$ is a product of many i.i.d. random variables, and we can use the central limit theorem in logspace: $f_{L}(s)$ becomes a lognormal distribution. Let
$t_{i}$ $=\ln r_{i}=\frac{s_{i+1}}{s_{i}},$ (S.101)
$\ln s$ $=\sum_{i=1}^{L-1}t_{i},$ (S.102)
then $t_{i}$ is a log-transformed Beta random variable with mean and variance
$\mu$ $=-\sum_{n=1}^{K-1}\frac{1}{n}=-H_{K-1},$ (S.103)
$\sigma^{2}$ $=\sum_{n=1}^{K-1}\frac{1}{n^{2}}=H^{(2)}_{K-1}$, (S.104)
Then for the sum of (L1)(L-1) such random variables, the distribution is
$f_{L}(s)$ $\simeq\frac{1}{\sqrt{2\pi}\sigma_{L}s}\exp\left\{\frac{(\ln s-\mu_{L})^{2}}{2\sigma_{L}^{2}}\right\},$ (S.105)
$\mu_{L}$ $=-(L-1)H_{K-1},$ (S.106)
$\sigma^{2}_{L}$ $=(L-1)H^{(2)}_{K-1}$. (S.107)
V.1 Sanity check: Compare lognormal with the exact solution for K=2
Let’s check the special case of $K=2$, which admits exact solution (Sec. II.2)
$f_{L}(s)=\frac{(-\ln s)^{L-2}}{(L-2)!}.$ (S.108)
Let $x=-\ln s$, also assume $L=L-2\gg 1$, then (with Jacobian factor $e^{-x}$)
$g_{L}(x)=f_{L}(s)e^{-x}=\frac{1}{L!}e^{L\ln x-x}$ (S.109)
Define $x:=L+\delta$, where we assume $\delta\sim O(\sqrt{L})$, we can rewrite
$\ln x$ $=\ln(\delta+L)$ (S.110)
$=\ln L+\ln(1+\frac{\delta}{L})$ (S.111)
$=\ln L+\frac{\delta}{L}-\frac{\delta^{2}}{2L^{2}}+O(L^{-3})$ (S.112)
Then
$\ln g_{L}(x)$ $=-\ln L!+L\ln L+\delta-\frac{\delta^{2}}{2L}-x$ (S.113)
$=L+\delta-\frac{\delta^{2}}{2L}-x-\frac{1}{2}\ln(2\pi L)$ (S.114)
$=-\frac{\delta^{2}}{2L}-\frac{1}{2}\ln(2\pi L),$ (S.115)
where in the second line we have used the Stirling approximation
$\ln L!=L\ln L-L+\frac{1}{2}\ln(2\pi L)+O(\frac{1}{L}).$ (S.116)
Therefore,
$g_{L}(x)=\frac{1}{\sqrt{2\pi L}}\exp\left\{-\frac{(x-L)^{2}}{2L}\right\}.$ (S.117)
Going back to $s=e^{-x}$, we have (for $K=2$)
$f_{L}(s)$ $=\frac{1}{\sqrt{2\pi(L-2)}s}\exp\left\{-\frac{[\ln s+(L-2)]^{2}}{2(L-2)}\right\},$ (S.118)
which for $L\gg 1$ becomes identical to the $K=2$ case in Eq.(S.105), as $\mu_{L}=-(L-1)H_{1}=-(L-1)$ and $\sigma_{L}^{2}=(L-1)H^{(2)}_{1}=(L-1)$.
V.2 RG flow of non-gaussianity
The normalized chunk size variable $s$ becomes a lognormal random variable for $L\gg 1$, we define the rescaled $O(1)$ random variable
$x=\frac{\ln s-\mu_{L}}{\sigma_{L}}.$ (S.120)
Let’s consider the effect of adding one level, $L\to L+1$.
$s_{L+1}=r_{L}s_{L},\;\;r_{L}\sim p_{B}=\text{Beta}(1,K-1),$ (S.121)
So using the definitions in Eq. (S.103) and Eq. (S.120), we can write
$x_{L+1}$ $=\frac{\ln s_{L+1}-L\mu}{\sqrt{L}\sigma}$ (S.122)
$=\sqrt{\frac{L-1}{L}}x_{L}+\frac{1}{\sqrt{L}}\frac{\ln r_{L}-\mu}{\sigma}$ (S.123)
$:=a_{L}x+b_{L}y,$ (S.124)
where we have isolated an $O(1)$ and an $O(L^{-1/2})$ contribution
$a_{L}$ $=1-\frac{1}{2D}+O(L^{-2}),$ (S.125)
$b_{L}$ $=\frac{1}{\sqrt{L}}.$ (S.126)
The log-variables are distributed as
$x$ $\sim\sigma\sqrt{L-1}sf_{L}(s):=P_{L}(x),$ (S.127)
$y$ $\sim\sigma sf_{2}(s):=P_{2}(y),$
where $P_{2}(y)$ is the pdf for the 1-step log-beta variable $y=\ln r_{L},\;r_{L}\sim p_{B}$, and $P_{L}(x)$ is the pdf of the (L1)(L-1) step log-variable in Eq. (S.120).
Therefore, the $L$ step log-variable (Eq. (S.122)) has pdf
$P_{L+1}(z)$ $=\int_{-\infty}^{\infty}P_{L}(x)P_{2}(y)\delta(z-a_{L}x-b_{L}y)dxdy$ (S.128)
$=\frac{1}{a_{L}}\int_{-\infty}^{\infty}P_{L}\left(\frac{z-b_{L}x}{a_{L}}\right)P_{2}(x)dx$ (S.129)
Fourier transforming Eq. (S.129) we obtain
$\hat{P}_{L+1}(k)=\hat{P}_{L}(a_{L}k)\hat{P}_{2}(b_{L}k).$ (S.130)
Eq. (S.129) is the “real-space” RG map transforming the pdf from $L$ to $L+1$, while Eq. (S.130) is the Fourier-space RG map.
By Central Limit Theorem (CLT), $P_{L}(z)$ becomes a normal distribution for $L\gg 1$, so let’s try to expand around this Gaussian fixed point,
$P_{L}(z)$ $=\phi(z)\left[1+\sum_{m\geq 0}^{\infty}\frac{g_{m}(L)}{m!}H_{m}(z)\right],$ (S.131)
$\phi(z)$ $=\mathcal{N}(0,1),$ (S.132)
where $g_{m}(L)$ are the coefficients of the non-gaussianity modes (not all modes are free to vary, see below), and $H_{m}(z)$ are the Hermite polynomials satisfying the following properties:
$H_{m}(z)$ $:=(-1)^{m}e^{\frac{z^{2}}{2}}\partial_{z}^{m}e^{-\frac{z^{2}}{2}},$ (S.133)
$H_{m}(z)\phi(z)$ $=(-1)^{m}\partial_{z}^{m}\phi(z),$ (S.134)
$n!\delta_{mn}$ $=\int_{-\infty}^{\infty}dz\phi(z)H_{m}(z)H_{n}(z).$ (S.135)
Now we impose the constraints that $P_{L}(z)$ needs to be normalized, and has zero mean, unit variance:
$1$ $=\int_{-\infty}^{\infty}dzP_{L}(z)=\int_{-\infty}^{\infty}dzP_{L}(z)H_{0}(z)$ (S.136)
$0$ $=\int_{-\infty}^{\infty}dzP_{L}(z)z=\int_{-\infty}^{\infty}dzP_{L}(z)H_{1}(z)$ (S.137)
$1$ $=\int_{-\infty}^{\infty}dzP_{L}(z)z^{2}=\int_{-\infty}^{\infty}dzP_{L}(z)[1+H_{2}(z)],$ (S.138)
where we have used the fact that the first three Hermite polynomials are:
$H_{0}(z)$ $=1$ (S.139)
$H_{1}(z)$ $=z$ (S.140)
$H_{2}(z)$ $=z^{2}-1.$ (S.141)
Using the orthogonality condition of the Hermite polynomials (Eq. (S.135)), we have $g_{0}=g_{1}=g_{2}=0$.
Next, since
$\mathcal{F}[\phi](k)$ $:=\hat{\phi}(k)=e^{-\frac{k^{2}}{2}},$ (S.142)
$\mathcal{F}[H_{m}\phi]$ $=(-1)^{m}\mathcal{F}[\partial_{z}^{m}\phi(z)]=(ik)^{m}\hat{\phi}(k),$ (S.143)
The Fourier transformed pdf becomes
$\hat{P}_{L}(k)=e^{-\frac{k^{2}}{2}}\left[1+\sum_{m\geq 3}g_{m}\frac{(ik)^{m}}{m!}\right].$ (S.144)
Let’s define
$\Phi_{L}(k)$ $=\ln\hat{P}_{L}(k)=\ln\mathbb{E}\left[e^{ikz}\right]$ (S.145)
$=-\frac{k^{2}}{2}+\ln\left[1+\sum_{m\geq 3}g_{m}\frac{(ik)^{m}}{m!}\right]$ (S.146)
$\approx-\frac{k^{2}}{2}+\sum_{m\geq 3}g_{m}(L)\frac{(ik)^{m}}{m!},$ (S.147)
then it becomes apparent that $g_{m}(L)$ are the cumulants of $P_{L}(z)$.
Now the goal is to express both sides of Eq. (S.130) in terms of these cumulant expansions, and match the coefficients. The RHS is:
$\Phi_{L}(a_{L}k)$ $=-\frac{1}{2}a_{L}^{2}k^{2}+\sum_{m}g_{m}(L)a_{L}^{m}\frac{(ik)^{m}}{m!},$ (S.148)
$\Phi_{2}(b_{L}k)$ $=-\frac{1}{2}b_{L}^{2}k^{2}+\sum_{m}\gamma_{m}(L)b_{L}^{m}\frac{(ik)^{m}}{m!}.$ (S.149)
Therefore (noting that $a_{L}^{2}+b_{L}^{2}=1$)
$\Phi_{L+1}(k)$ $=-\frac{1}{2}k^{2}+\sum_{m}\left[a_{L}^{m}g_{m}(L)+b_{L}^{m}\gamma_{m}(L)\right]\frac{(ik)^{m}}{m!}$ (S.150)
$=-\frac{1}{2}k^{2}+\sum_{m}g_{m}(L+1)\frac{(ik)^{m}}{m!},$ (S.151)
we can read off the RG update of the coefficients:
$g_{m}(L+1)=a_{L}^{m}g_{m}(L)+b_{L}^{m}\gamma_{m}(L).$ (S.152)
Next, we relate $g_{m}$ with $\gamma_{m}$, using the fact that they come from the same underlying family of distributions $f_{L}(s)$ (Eq. (S.127)). Let’s measure everything in terms of the cumulant of the unscaled zero-mean log-variable $(\ln r_{i}-\mu)$, $\kappa_{m}$:
$\gamma_{m}[y]=\gamma_{m}\left[\frac{\ln r_{i}-\mu}{\sigma}\right]=\frac{\kappa_{m}}{\sigma^{m}},$ (S.153)
whereas
$g_{m}[x]$ $=g_{m}\left[\frac{\sum_{i}^{L-1}\ln r_{i}-(L-1)\mu}{\sqrt{L-1}\sigma}\right]$ (S.154)
$=\frac{(L-1)\kappa_{m}}{(L-1)^{m/2}\sigma^{m}}$ (S.155)
$=\frac{\gamma_{m}}{(L-1)^{\frac{m-2}{2}}}.$ (S.156)
Therefore, inserting Eq. (S.156) back to Eq. (S.152) we have
$g_{m}(L+1)$ $=g_{m}(L)\left[a_{L}^{m}+b_{L}^{m}(L-1)^{\frac{m}{2}-1}\right]$ (S.157)
$=g_{m}(L)\left[1-\frac{m-2}{2D}+O(L^{-2})\right].$ (S.158)
We can turn Eq. (S.158) into a differential equation
$\frac{d\ln g_{m}(L)}{dD}$ $=-\frac{m-2}{2}\frac{1}{L}.$ (S.159)
Solving Eq. (S.159) gives
$g_{m}(L)$ $=L^{-\Delta_{m}},$ (S.160)
$\Delta_{m}$ $=\frac{m-2}{2},\;\;m\geq 3,$ (S.161)
where $\Delta_{m}$ are the scaling dimensions of the non-gaussianity fluctuation modes. Following the RG conventions, the beta-function of the RG flow is (defining $\ell=\ln L$)
$\frac{dg_{m}(\ell)}{d\ell}$ $=-\frac{m-2}{2}g_{m}(\ell).$ (S.162)
Since $g_{m}(L)$ are the cumulants of $P_{L}$ (Eq. (S.147)), Eq. (S.160) confirms that the CLT we have assumed is self-consistent: all $m\geq 3$ cumulants are irrelavant directions of the RG flow, and all these higher cumulants vanishes asymptotically as $L\to\infty$.
This derivation can be thought of as a RG derivation of the central limit theorem, in a spirit similar to the treatment in [59] for continuous RG transformations, but with important distinction that this applies to discrete random $K$-ary trees.